OpenStudy (anonymous):

Elena's bus runs every 20 minutes. If she arrives at her bus stop at a random time, what is the probability that she will have to wait at least 5 minutes for the bus if it is running on schedule? A)5% B)20% C)50% D)75%

OpenStudy (anonymous):

5/20 = 0.25 (25%) 100 - 25 = 75% D, 75%

OpenStudy (anonymous):

thank you